OpenStudy (sneville):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16) A. y = x + 6 B. y = x – 6 C. y = x + 14 D. y = x

OpenStudy (will.h):

First of all parallel equation share the same slope so our slope for the new equation would be 1 and we have the point (2,16) We can use slope intercept form to make up the equation Y-y1 = m(x - x1) Where m is the slope Y - 16 = 1(x -2) Y - 16 = x -2 Y = x + 14 Hope that helps. If you have questions about it ask me